OpenStudy (anonymous):

Of 18 students in a class, 10 are girls. The teacher selects a student at random. What is P(not a girl)? A. 2/3 B. 5/9 C. 4/9 D. 1/3

OpenStudy (anonymous):

Medal and fan if help

OpenStudy (btaylor):

P(girl) = 10/18 = 5/9 P (not a girl) = 1-P(girl)=1-5/9 = 4/9